Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day13


Crew 100A Log Book for February 24, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA:

EVA #19 (11:02-11:37),

EVA#20(12:03.12:35),

EVA#21(16:16-16:44),

EVA#22(16:54-17:17)

Brief Narrative of Field Mission Results:

During the morning we did two EVAs with no special communication protocol to serve as calibration data for all the other EVAs. They were done on ATV to locations no more than 3 km distant from the Habitat.

After lunch we had time for the debriefing of the habitability experiment on which all the crewmembers exposed their view on the MDRS experience and mentioned issues that needed to be improved when regarding long isolation periods like those expected for long duration space flights.

After that, there was time for the two last EVAs of the stress experiment, which included gathering geological samples and fixing power connections, always with unexpected emergencies.

Hopefully tomorrow, between cleaning periods and development of the range extender for the parrot, we will be able to visit a bit of the sites nearby.

Engineering/Hab Maintenance: The system of the GreenHab is slowly getting into place.

Plans for Tomorrow:  Wrap up of the EVAs to test the effect of communication on stress during an EVA. Report all the work done. Clean the Habitat. EVAs with ATVs to check the routes near the Hab that we did not have the chance to visit.

Reports scheduled for today: Engineering, Commander, Journalist, Science, EVA #19, EVA#20, EVA#21 and EVA#22.

Inventory: N/A

Support Requested: The pump from the trailer to the Hab tank doesn't seem to be working. For our crew it will be all right with the quantity of water but there might be a problem for the new crew.

Miscellaneous: The GPS unit at the Hab seems to be worse and worse.

Sometimes it does not inform on the directionality when within 100m of the target. Hopefully the new one that is being shipped will solve this problem.


Commander's Report

Ivo Ferreira Reporting

Fake emergencies

If there is something that we are getting used here in Mars is to have a lot of fake emergencies.

One of the experiments was designed to create stressful situations induced by communication problems, the idea is to study the impact of different communication protocols on the stress of the EVA. We have had so many electrical break downs, contaminated samples and lack of oxygen that it is a miracle we are still alive.

Today, during the morning and the afternoon, we completed 4 EVAs as part of this experiment. During two of them we managed to achieve the scientific goals and prevent certain death of our crew members but on the other 2 we were not so successful. Ask Ayako and Aydrey and they might tell you that it is not a nice thing to run out of oxygen here in Mars.

We also managed to have a nice discussion about the habitability issues of our experience here in Mars during the debriefing of Ayako's experiment. It was our time to complain about the quality of the spacesuits, electrical noises, toilet, showers, privacy (or lack of it) and, most importantly, the FOOD. As Kévin mentioned: "good food is the key of a good mood".

Hopefully some of these issues can be changed before our 12 months trip back to Earth that will start 2 days from now.

Tomorrow, besides spending some time with the parrot and Max, it will be the wrapping-up-cleaning-sightseeing day. We will try to take some pictures of this beautiful planet to share with you guys back on Earth.


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.75

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.48

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 5 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 16,5

· Trailer Potable Water Tank Level (inches from bottom): 3/4Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading: 424946

Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little. The process is slowly getting into place.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): no

· Oil Checked (yes/no): No

· Tire Status: checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status: filled

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 68

· Thermostat setting downstairs (Farenheit): 60

· Actual temperature downstairs (Farenheit): 57


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Science Report

Ayako Ono Reporting

Habitability Debriefing (Workshop)

Firstly, we discovered problems on IVA (all in the Hab) and EVA about following factors: Social, Environmental-Physiological, Psychological, Operational.

Secondly, we read each factor’s topics.

Thirdly, we discussed about what we found as problems and solutions and additional comments for future long-term missions as others. E.g. We found the importance of good foods, radio, GPS, and EVA. Too many tasks and bathroom breaks are problems. Joking, movies, music, painting, writing, singing, and sports (Wii: game) will be helpful as solutions.


EVA 19 Report

Kévin Théatre Reporting

EVA #19, 11:00-11:35

Crew Members: Ivo Ferreira, Romain Benchenafi, Kévin Théatre

Site Location:

N38.42980° W110.78764° - N38.41330°-W110.78191°

Transit Mode: ATVs

Objectives: Mission 6 (measure the stress induced by radio-communications on EVA missions).

Results: the 3 critical objectives were a complete success.

Lessons Learned: None


EVA 20 Report

Ayako Ono Reporting
EVA #20, 12:03-12:35
Crew Members: Ivo Ferreira, Audrey Lan San and Ayako Ono
Site Location:
N38.42980° W110.78764° - N38.41330°-W110.78191°
Transit Mode: ATVs
Objectives: Mission 6 (measure the stress induced by radio-communications on EVA missions).
Results: the 1 critical objective was completed, then we needed to go back.
Lessons Learned: Selection of road will cut the time.

EVA 21 Report

Romain Benchenafi Reporting

EVA #21, 16:23-16:44

Crew Members: Benchenafi Romain, Kévin Théatre

Site Location:

N38.39360 W110.79034, N38.39048 W110.78458

Transit Mode: ATVs

Objectives: study communication related to stress

Results: Managed to comply with all objectives

Lessons Learned: None


EVA 22 Report

Romain Benchenafi Reporting

EVA #22, 16:54-17:17

Crew Members: Benchenafi Romain, Kévin Théatre

Site Location: N38.39360 W110.79034, N38.40655 W110.79134

Transit Mode: ATVs

Objectives: study communication related to stress

Results: Didn't comply with the emergency

Lessons Learned: The GPS is less accurate when it is running out of batteries


Chef's Report

Ayako Ono Reporting

1a) Today was:

 a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No, it was not.

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each

-       Santa Fe Black Beans

-       Sliced Potatoes with Cheese Sauce

5)  Lessons learned  (if any)

No

6) Comments/questions for Kim and Jean, the food study investigators - any other thoughts or suggestions?

Good foods change crew members’ mood.

Cooking is something creative and if would be better if astronauts can cook with some dried foods.

I didn’t understand why we need to write recipe when we became a cooker. Also, food diary was too hard to describe all food’s materials. If you could reduce questions, would be much better, like five short questions.


Journalist Report

Romain Benchenafi Reporting

A busy day, again for the crew 100A, as we wanted to finish all our experiments.

Indeed we wished to do all the cleaning of the Hab tomorrow and conclude our mission.

Unfortunately we still have some work to achieve during what will be our last full day on Mars.

It ends soon!